OpenStudy (anonymous):

Debra earns $7.50 an hour, plus tips, as a waitress. Suppose she works 24 hours in one week. How much money must she receive in tips to earm more than $250 that week?

OpenStudy (anonymous):

let x be how much she earns in a week let y be her tips x=(wagerate)(no. of hours) x=(7.50)(24) x=180 x+y= 250 dollars 180+y=250 y=70 Debra must get more than 70 $ worth of tips.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."